OpenStudy (anonymous):

how to plug in trigonometry problems into a TI-83 calculator. Ex: sin X = (11/26)

OpenStudy (anonymous):

well i got 2.36 repet ok

OpenStudy (anonymous):

know x is 2.36

OpenStudy (anonymous):

man ok hold on

OpenStudy (anonymous):

wait its sin X = (11/26) my teacher key sheet said its 26?

OpenStudy (anonymous):

i mean sin X = (11/25)

OpenStudy (anonymous):

I just need help plugging it to a calcutlator

OpenStudy (anonymous):

calculator*

OpenStudy (anonymous):

sine is 26

OpenStudy (anonymous):

do u know how

OpenStudy (anonymous):

tho

OpenStudy (anonymous):

no i dont know how to plug it into a calculator, thats the only part i need help on. I have a TI-83

OpenStudy (anonymous):

26x=(11/26)

OpenStudy (anonymous):

try that

OpenStudy (anonymous):

then tell me what happen

OpenStudy (anonymous):

woah i got 0?

OpenStudy (anonymous):

i just need help from a questions that needs to plugged in but i dont know how, from sin X = (11/25) how do i plug it in? from there it should be 26

OpenStudy (anonymous):

wee the x is put the 26

OpenStudy (anonymous):

lol srry i mean were

OpenStudy (anonymous):

what is a TI-83?is that some special kind of calculator?...if it has shift all you have to do is press shift then sin(11/25)...

OpenStudy (anonymous):

no that is the problem

OpenStudy (anonymous):

what do you mean dkglenn?

OpenStudy (anonymous):

ti-83 that is what he hase know ok but he need the sin and the sin is 26 but he need to plug it in the calculater.

OpenStudy (anonymous):

that he does not know how

OpenStudy (anonymous):

did u get it temmayy

OpenStudy (anonymous):

i think he did lol

OpenStudy (anonymous):

nah, im still trying to plug in the numbers into my calculator to end up with 26, cause the answer sheet says 26 but everything i tryed it still hasnt worked, the problem is sin X (11/25) int he end X will be 26

OpenStudy (anonymous):

does this calculator have a sign button?

OpenStudy (anonymous):

i meant 'sin'

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then you must hit the yellow button that says 2nd before you hit the sine button...2nd is equivalent to shift...did you try that?

OpenStudy (anonymous):

OHH SNAP THANKS MAN!

OpenStudy (anonymous):

glad that it worked...

OpenStudy (anonymous):

God_in_me is correct - he is showing you how to access the sin^(-1) function. sin^(-1) is the inverse function of the sin function. It is also know as the Cosecant. So to solve your equation this is what your actually doing. "Solve for x... \[\frac{1}{\sin}*\sin(X)=\frac{1}{\sin}*\frac{11}{26}=(\sin^{-1})(11\div26)\]

OpenStudy (anonymous):

It's kind of like dividing both sides by 5 in this equation. \[5x=100\] But you need to know how to find the inverses of your three main trig functions.